OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of f on the given interval. f(x)=x^3-9x^2; [0,10]

OpenStudy (helder_edwin):

is this calculus? if so, find the first derivative

OpenStudy (anonymous):

yes this calculus

OpenStudy (helder_edwin):

so find the first derivative

OpenStudy (helder_edwin):

can u do it?

OpenStudy (anonymous):

not really it takes me a mintue

OpenStudy (anonymous):

@vjydubey I cant read the outcome

OpenStudy (helder_edwin):

in this case u have to follow two simple rules \[ (u\pm v)'=u'\pm v' \] \[ (x^n)'=n\cdot x^{n-1} \]

OpenStudy (anonymous):

okay..follow steps 1) find critical points of curve and check whther they are local maxima or minima 2)evalute function at that point 3)also evalut them at boundries for maximums which are not maxima

OpenStudy (helder_edwin):

so, since \[ f(x)=x^3-9x^2 \] then \[ f'(x)=(x^3)'-(9x^2)'=3x^2-9(2x)=3x^2-18x \]

OpenStudy (anonymous):

@helder_edwin what numbers go where... im confuse

OpenStudy (anonymous):

okay okay im following you

OpenStudy (helder_edwin):

sorry, this web site is acting weird

OpenStudy (helder_edwin):

at least from my end.

OpenStudy (helder_edwin):

once you have the first derivative, you solve the equation \[ y'=0 \]

OpenStudy (helder_edwin):

so now we solve \[ 0=f'(x)=3x^2-18x=3x(x-6) \] from which we have that x=0 or x=6

OpenStudy (helder_edwin):

both solutions lie in the interval given

OpenStudy (helder_edwin):

got it, so far?

OpenStudy (anonymous):

yes I solved it

OpenStudy (helder_edwin):

great